Ex. 4 A hole of cross-sectional area 100 cm^2 is cut in a piece of steel at 20 degrees C. What is the area of the hole if the steel is heated from 20 degrees C to 100 degrees C?

Ex5. A beaker is filled with 2000 cm^3 of turpentine at room temp. of 20 degree C. If beaker and contents are put in a refrigarator so that the temp. becomes 5 degrees C, what is the new volume of the turpentine? For Turpentine β= (900 x e^-6)/Degree C. Ignore the contraction of the beaker.

Ex 8 A volume if 22 m^3 of nitrogen gas at 20 degrees C is heated under constant pressure to 167 degrees C. What is the new volume?

Ex.9 20 l. of gas are kept under a pressure of 1 atm. at a temperature of 273 K . The gas temperature is lowered to 90 K. The pressure is increased to 1.5 atm. What is the new volume?
Answer #1

Ex-4

Superficial or areal Coefficient of expansion of steel is

2×7.2×10^-6 /°C = 1.44 ×10^-5/°C

So new area A = 100(1+1.44×10^-5×80)=100×(1.001152)

Or A = 100.1152 cm2

Ex-5

New volume V = 2000(1+ 900×10^-6×(5-20) = 1973 cm^3

Ex-8

From gas law V1/T1= V2/T2

So V2 = 22×(273+167)/(273+20) = 33.037 m3

Ex-9

As P1V1/T1= P2V2/T2

SO V2= (20×1/273)×90/1.5 = 4.395 l